Selma needs to prepare 180 pounds of coffee selling for $4.72 per pound. She plans to do this by blending together a high-quality bean costing $5.50 per pound and a cheaper pound at $3.50 per pound. To the nearest pound find out how much high quality beans and how many of the cheaper beans coffee beans should be blended. She should blend _____ lbs of the higher costing bean.
She should blend ______ lbs of the cheaper costing bean.

Answer by stanbon(75887)

Selma needs to prepare 180 pounds of coffee selling for $4.72 per pound. She plans to do this by blending together a high-quality bean costing $5.50 per pound and a cheaper bean at $3.50 per pound. To the nearest pound find out how much high quality beans and how many of the cheaper beans coffee beans should be blended.
--------
Equations:
Quantity Eq:: t + f = 180 lbs
Value Eq::: 3.50t+5.50f=4.72*180
--------------------------------------
Multiply thru the Quantity Eq by 350.
Multiply thru the Value Eq by 100.
---------------------------------------
350t + 350f = 350*180
350t + 550f = 472*180
----
Subtract and solve for "f":
200f = 122*180
f = (9/10)122
f = 109.8 lbs (amt. of higher priced coffee needed)
---
Solve for "t":
t + f = 180
t = 180-109.8 = 70.2 lbs (amt. of lower priced coffee needed)
=======================================
Cheers,
Stan H.
=======================================
She should blend _____ lbs of the higher costing bean.
She should blend ______ lbs of the cheaper costing bean.